OpenStudy (anonymous):

4 pencils and 5 pens cost $2.00. 3 pencils and 4 pens cost $1.58. find the cost of a pencil and the cost of a pen

OpenStudy (anonymous):

using the same as the last ?

OpenStudy (anonymous):

let cost of each pencil iy x$ and that of each pen is Y$ 4x+5y=2 3x+4y=1.58 solve this two equations.....u will get x and y

OpenStudy (anonymous):

from 1st equation x=(2-5y)/4 putting it in the second equation... 3(2-5y)/4+4y=1.58 6-15y+16y=6.32 y=.32$ and x=(2-1.6)/4=.1

OpenStudy (anonymous):

i think y=.32 not .032

OpenStudy (anonymous):

yes yes mistake, it is .32, sorry. and thanks dipank, y=8/25 = 0.32

OpenStudy (anonymous):

what about the previous question chloebella

OpenStudy (anonymous):

together a skateboard and a bicycle cost $199.00. the skateboard costs $51.00 less than the bicycle. how much does each cost?????

OpenStudy (anonymous):

let skate..is x and bicycle is x+51$ so x+x+51=199 2x=148 x=74 bicycle=125

OpenStudy (anonymous):

2 cars leave town at the same time going in the same direction. 1 travels 55mi/h and the other travels 48mi/h. in how many hours will they be 49 miles aprt?????

OpenStudy (anonymous):

in one hour they are apart (55-48)=7 mile so they will be apart 49 mile in 49/7=7 hour

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

if u have asked this questions separately then i could have got 4 medals ....

OpenStudy (anonymous):

sorry i will do that